Let w = 2 + 5i and z = -3 + 4i.
Compute 2ww + |z|^2 in Cartesian coordinates. The second w is the conjugate of w, my apologies, I don't know how to show it.
If I understand correctly I'm to get an x and y coordinate however, when I try to compute I end up with a single number. Any help would be greatly appreciated.

Answer by Alan3354(69443)

Let w = 2 + 5i and z = -3 + 4i.
Compute 2ww + |z|^2 in Cartesian coordinates.
------------------
2(2 + 5i)*(2 - 5i) = 2*(4+25) = 58
|z| = the magnitude of z, = sqrt(3^2 + 4^2) = 5
--> 63 or (63,0) since there's no i term.